Q1)MR. ANG INVESTED $865 IN A BUSINESS MAKING GIFT BASKETS. HE MAKES $13.28 FOR EACH GIFT BASKET THAT HE SELLS. WRITE A FUNCTION THAT REPRESENTS THE SITUATION. HOW MANY GIFT BASKETS MUST MR. ANG SELL UNTIL HE MAKES A PROFIT?'
Q2) The number of bacteria cells in a culture can be modelled by the function B(t)= -0.03(t-4)^2+12
1
Expert's answer
2020-11-17T11:35:18-0500

"Q1) P(x)=R(x)-C(x)" Given "C(x)=\$865, R(x)=\$13.28\cdot x"

Then


"P(x)=13.28x-865"

If "P(x)\geq0"


"13.28x-865\geq0"

"x\geq\dfrac{865}{13.28}"

"x\geq66"

"66^{th}" basket will give a profit.

Q2) The question was not described.